OpenStudy (anonymous):

s=180n-360 solve for n

OpenStudy (amistre64):

s+360 ------ = n 180

OpenStudy (anonymous):

n=s+360/180

OpenStudy (amistre64):

uzma is notationally wrong lol

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

n = (s + 360) / 180

OpenStudy (anonymous):

n=(s+360)/180 When s=1, n=361/180=2.0055 When s=2, n=362/180=2.0111

OpenStudy (anonymous):

whos right haha